# Vortices and Factorization

### v2

@@ 12 Time-Dependent Darboux Transformations, Calogero-Moser Systems and Locus Configurations
$l$ conditions for the first and $m$ conditions for the second specie respectively. It turned out [2] [^17] Conditions (98) first appeared in this famous paper by Airault, McKean and Moser, who introduced the term “locus”. that set of solutions of (98) is not empty only if the number of particles the locus is formed of is a triangular number $n(n+1)/2$. Then, it was shown [1] that for this number, roots of the Adler-Moser polynomials $P_{n}$ constitute complete solution of (98). As we saw in section 3, in systems of vortices in static equilibrium $l=n(n-1)/2$ and $m=n(n+1)/2$ particles form two loci that are roots of two consecutive Adler-Moser polynomials $P_{n-1}(z,t_{1},\dots,t_{n-2})$ and $P_{n}(z,t_{1},\dots,t_{n-1})$.
Conditions (98) also must hold for each specie of translating configurations (5) with ${\cal Q}_{i}=\pm 1$. We recall that for translating configurations $\deg p=\deg q$, i.e. $l=m$ (see section 8). This means that when $k\not=0$, $p$ and $q$ are Adler-Moser polynomials of the same degree, but with different sets of parameters: If $p=P_{n}(z,t_{1},\dots,t_{n-1})$, then $q=P_{n}(z+s_{0},s_{1},\dots,s_{n-1})$, where $s_{0},\dots,s_{n-1}$ are functions of $t_{1},\dots,t_{n-1}$ and $k$.
We also recall that, according to section 8, the rational Baker-Akhieser function $\psi=C(k)(q/p)e^{kz}$ determines translating configuration. It is an eigenfunction of a Schrodinger operator that is obtained by a sequence of Darboux transformations from free Schrodiger operator, i.e. $H\psi=k^{2}\psi$, where $H=-\partial_{z}^{2}-2(\log p)^{\prime\prime}$. From this we get the bilinear relation for $p$ and $q$:
$p^{\prime\prime}q-2p^{\prime}q^{\prime}+pq^{\prime\prime}+2k(q^{\prime}p-p^{\prime}q)=0.$ (99)
This relation coincides with the Tkachenko equation for translating configurations. Set of its solutions contains all translating patterns with ${\cal Q}_{i}=\pm 1$. The Baker-Akhieser function exists for any locus of $p$ and is unique for a given locus (see e.g. [13, 12] and references therein). Since the locus conditions are necessary conditions of an equilibrium, we come to the conclusion that all possible translating configurations consisting of a finite number of vortices with ${\cal Q}_{i}=\pm 1$ are determined by the Baker-Akhieser functions and are given by (69).
In the trigonometric (periodic) case, for each of two species of opposite charges the locus conditions write as
@@
It is known (see e.g. [13, 12] and references therein) that all trigonometric locus configurations are given by roots of soliton $\tau$-functions of the KdV hierarchy. Therefore any pair $p$ and $q$ that does not have common roots and solves the periodic Tkachenko equation
@@
consists of soliton $\tau$-functions. These functions are obtained with the help of factorization chain (34) starting from the free Schrodinger operator. In this chain of transformations
$H_{n}\psi^{(n)}=\eta_{n+1}\psi^{(n)},\quad H_{n}=-\partial_{z}^{2}-2(\log\tau_{n})^{\prime\prime},\quad\psi^{(n)}=\frac{\tau_{n+1}}{\tau_{n}},\quad\tau_{0}=1,$
i.e.
$\left(-\partial_{z}^{2}-2(\log\tau_{n})^{\prime\prime}\right)\left[\frac{\tau_{n+1}}{\tau_{n}}\right]=\eta_{n+1}\frac{\tau_{n+1}}{\tau_{n}}.$ (101)
Since degree of the leading term of the Laurent expansion of $\tau_{n+1}/\tau_{n}$ in $\exp({\rm i}z)$ equals $d_{n+1}-d_{n}$, where $d_{n}=\deg\tau_{n}$, from (101) we conclude that $\eta_{n+1}=(d_{n+1}-d_{n})^{2}$. Then, rewriting (101) in the bi-linear form, we get the Darboux chain for $\tau_{n}$:
@@
The periodic analog of the Tkachenko equation (100) is of the same form as the Darboux chain equation (102). Then, taking the above fact and the fact that $p$ and $q$ are necessarily soliton $\tau$-functions (as follows from locus conditions) into account, we come to the conclusion that $p$ and $q$ are soliton $\tau$-functions related by Darboux transformation. Therefore, all possible static street patterns consisting of vortices of two species of opposite strengths are given by (78,79).
For translating streets we use arguments similar to those of the rational case and come to the conclusion that the complete set of translating street patterns with ${\cal Q}_{i}=\pm 1$ is determined by (78,82). It is worth mentioning that from the soliton theory it follows that here the relationship between $p$ and $q$ holds
@@
where $\tau_{n}$ is given by (78,79). Without going into much details, we note that the above relationship can be derived using the Sato formula for the Baker-Akhieser function of the KP hierarchy and the correspondence between Hirota and Wronskian representations of the $n$-soliton $\tau$-functions (for introduction to theory of solitons see e.g. [39]).
Closing this section, we note that, in contrast to the cases considered above, the $\Lambda=2$ case is not exceptional as no decoupling into sub-systems of different species takes place there.
@@ 13 Conclusions and Open Problems
In the present paper we reviewed developments in application of factorization methods to the theory of point vortex patterns that took place during several last decades. In the $\Lambda=1$ case, these have connections with the theory of KdV hierarchy of integrable PDEs and the bi-spectral problem. In the $\Lambda=2$ case, connections with integrable hierarchies have been found by Demina and Kudryashov [15] and further developed by us in [28]. In difference from the KdV hierarchy, the Lax operators of these hierarchies are differential operators of the third order  [^18] We recall that in $\Lambda=1$ case the Lax operator is of the second order: Lax operator of the KdV hierarchy is a Schrodinger operator.. Indeed, in the $\Lambda=2$ case the bilinear equation (8) for $q_{n}$ and $p_{n}$ can be rewritten in the form of the Schrodinger equation at zero energy level
$\left(-\partial_{z}^{2}+u\right)\phi=0,\quad u=-6(\log q_{n})^{\prime\prime},$ (103)
where $\phi=p_{n}/q_{n}^{2}$ or $\phi=p_{n-1}/q_{n}^{2}$. According to (24), we can replace $p_{n}$ with $q_{n+1}^{\prime}q_{n}-q_{n+1}q_{n}^{\prime}$ and get the following third-order differential equation for $\psi(z)=q_{n\pm 1}(z)/q_{n}(z)$
$L\psi=0,\quad L=\partial_{z}^{3}-u\partial_{z}.$
The third-order operator $L$ is a Lax operator of the Sawada-Kotera hierarchy and $q_{n}$ is its polynomial $\tau$-function. Such polynomials can be constructed with help of the Darboux transformations at “zero level” [28]
[…]
where primitives include arbitrary constants of integration. Corresponding third-order operators can be obtained by applications of sequence of the above transformations to $\partial_{z}^{3}$, while $q_{n}$ can be expressed in form of a Pfaffian [20].
Here we just note that, since $L$ is not a Schrodinger operator, the Baker-Akhiezer function $\psi(k,z)=P_{n}(k,z)/q_{n}(z)e^{kz}$ of the Sawada-Kotera hierarchy
$L\psi(k,z)=k^{3}\psi(k,z)$
does not correspond to stationary configurations when $k\not=0$. Also, soliton solutions of this hierarchy do not correspond to stationary configurations on the cylinder. Only when $k=0$, the above eigenvalue problem reduces to solution of the Schrodinger equation (103) with $\phi(z)=\lim_{k\to 0}k^{N_{n}}\frac{\partial\psi(k,z)}{\partial z}$ corresponding to static configurations. Here factor $k^{N_{n}}$, with suitably chosen $N_{n}$, insures finiteness of the limit.
In difference from the $\Lambda=1$ case, the $\Lambda=2$ case seems not to have as many exceptional properties: In particular, the fact that the soliton Baker-Akhiezer functions of the Sawada-Kotera hierarchy are not eigenfunctions of the Schrodinger operators suggests absence of non-trivial non-terminating sequences on the cylinder in the $\Lambda=2$ case.
@@
We searched for sequences of configuration on the cylinder by direct solution of the trigonometric analog of the bilinear equation (8)
$\tau_{n-1}^{\prime\prime}\tau_{n}-2\gamma_{n}\tau_{n-1}^{\prime}\tau_{n}^{\prime}+\gamma_{n}^{2}\tau_{n-1}\tau_{n}^{\prime\prime}+(d_{n-1}-\gamma_{n}d_{n})^{2}\tau_{n-1}\tau_{n}=0,\quad\gamma_{n}\gamma_{n+1}=1,$ (104)
@@
and in the $\Lambda=2$ case we found only short terminating sequences, for example
$\tau_{0}=\xi-\frac{1}{\xi},\tau_{1}=\xi^{6}-4\xi^{4}+5\xi^{2}+\frac{5t_{1}}{\xi^{2}}-\frac{4t_{1}}{\xi^{4}}+\frac{t_{1}}{\xi^{6}},\,\tau_{2}=\xi^{5}-5\xi^{3}-t_{2}\xi+\frac{t_{2}}{\xi}-\frac{5t_{1}}{\xi^{3}}+\frac{t_{1}}{\xi^{5}},$
@@
where $\xi=e^{{\rm i}z}$ and $\gamma_{0}=2$. The above example can be continued one more step if we impose condition $t_{1}=t_{2}^{2}/5$. Then we will get another terminating sequence with $\tau_{1}=\tau_{1}(\xi,t_{2})$, $\tau_{2}=\tau_{2}(\xi,t_{2})$ and $\tau_{3}=\tau_{3}(\xi,t_{2},t_{3})$ , where $\deg\tau_{3}=14$.
@@
Without going into much detail we just note that in such type of sequences $\tau_{1}$ is a solution of an ODE that is reducible to the hypergeometric form. Then we can apply KWCC transformations (60) to $\psi_{i}=\tau_{i}^{\gamma_{i}}/\tau_{i-1}$, $i=1,2$, to continue the sequence. The transformation can be applied at most twice. Terminating configurations obtained in this way can depend on no more than two non-trivial free parameters (they depend on two parameters when $\deg\tau_{1}=6,8,10,\dots$). Here, a natural question arises: If the above procedure generates all possible static vortex streets patterns of two distinct species of vortices with $Q_{i}\in\{-1,2\}$
Regarding the static patterns formed by finite number of vortices, we may ask the question: Whether the main sequences (23), (25), (26) for the $\Lambda=1$ and $\Lambda=2$ cases are the only possible non-trivial non-terminating sequences of the KWCC transforms. Are there non-trivial sequences (60) with $\gamma_{i}$ different from (61,62) that do not terminate?
@@
Concluding this review, we would like to mention related problem from the theory of algebraically integrable systems, because our interest in vortex patterns stems from that theory: In the case of odd family of periodic configurations, it is convenient to rewrite the bi-linear chain (102) in terms of homogenous polynomials in two variables, replacing $\tau_{n}(z)=\prod_{i=1}^{d_{n}}\sin(z-z_{i})$ with
@@
In the new variables, the chain (102) takes the following form

### v3

@@ 14 Time-Dependent Darboux Transformations, Calogero-Moser Systems and Locus Configurations
with $l$ conditions for the first species and $m$ conditions for the second species respectively.
It turned out [2] [^32] Conditions (118) first appeared in this famous paper by Airault, McKean and Moser, who introduced the term “locus”. that set of solutions of (118) is not empty only if the number of particles the locus is formed of is a triangular number $n(n+1)/2$. Then, it was shown [1] that for this number, roots of the Adler-Moser polynomials $P_{n}$ constitute complete solution of (118). As we saw in section 3, in systems of vortices in static equilibrium, $l=n(n-1)/2$ and $m=n(n+1)/2$ particles form two loci that are roots of two consecutive Adler-Moser polynomials $P_{n-1}(z,s_{1},\dots,s_{n-2})$ and $P_{n}(z,s_{1},\dots,s_{n-1})$.
Conditions (118) must also hold for each species of translating configurations (5) with ${\cal Q}_{i}=\pm 1$. We recall that, due to neutrality condition (8), for translating configurations $\deg p=\deg q$, i.e., $l=m$. This means that when $k\not=0$, both $p$ and $q$ are Adler-Moser polynomials of the same degree, but with different sets of parameters. As a result, according to section 10, all possible translating configurations consisting of a finite number of vortices with ${\cal Q}_{i}=\pm 1$ are given by (91). Specifically, $q=P_{n}(z,s_{1},\dots,s_{n-1})$, $p=P_{n}(z+\tilde{s}_{0},\tilde{s}_{1},\dots,\tilde{s}_{n-1})$, where $\tilde{s}_{0},\dots,\tilde{s}_{n-1}$ are functions of $s_{1},\dots,s_{n-1}$ and $k$.
In the trigonometric (periodic) case, for each of two species of opposite charges the locus conditions are writen as
@@
It is known (see e.g. [17, 18]) that all trigonometric locus configurations are given by the roots of soliton $\tau$-functions of the KdV hierarchy. Therefore, any pair $p$ and $q$ that does not share common roots and satisfies the periodic Tkachenko equation
@@
consists of soliton $\tau$-functions. On the other hand, such $\tau$-functions determine potentials in the factorization chain (35), which begins with the free Schrodinger operator. In this chain of transformations
$H_{n}\psi^{(n)}=\eta_{n}\psi^{(n)},\quad H_{n}=-\partial_{z}^{2}-2(\log\tau_{n})^{\prime\prime},\quad\psi^{(n)}=\frac{\tau_{n+1}}{\tau_{n}},\quad\tau_{0}=1,$
i.e.,
$\left(-\partial_{z}^{2}-2(\log\tau_{n})^{\prime\prime}\right)\left[\frac{\tau_{n+1}}{\tau_{n}}\right]=\eta_{n}\frac{\tau_{n+1}}{\tau_{n}}.$ (120)
Since the degree of the leading term of the Laurent expansion of $\tau_{n+1}/\tau_{n}$ in $\exp({\rm i}z)$ equals $d_{n+1}-d_{n}$, where $d_{n}=\deg\tau_{n}$, we can conclude from (120) that $\eta_{n}=(d_{n+1}-d_{n})^{2}$. Then, rewriting (120) in the bi-linear form, we obtain the Darboux chain for $\tau_{n}$:
@@
The periodic analog of the Tkachenko equation (119) has the same form as the Darboux chain equation (121). Taking this into account, along with the fact that $p$ and $q$ are necessarily soliton $\tau$-functions (as follows from locus conditions), we conclude that $p$ and $q$ are soliton $\tau$-functions related by a Darboux transformation. Therefore, all possible static street patterns consisting of vortices of two species of opposite strengths are defined by (100,101).
For translating vortex streets, we use arguments similar to those of the rational case, which leads us to the conclusion that the complete set of translating street patterns with ${\cal Q}_{i}=\pm 1$ is determined by (100,104). It is worth mentioning that, according to the soliton theory, the following relationship between $p$ and $q$ holds
@@
where $\tau_{n}$ is given by (100,101). Without going into much detail, we note that this relationship can be derived using the Sato formula for the Baker-Akhieser function of the KP hierarchy, as well as the correspondence between Hirota and Wronskian representations of the $n$-soliton $\tau$-functions (for an introduction to the soliton theory, see e.g. [25, 49]).
@@ 15 Conclusions and Open Problems
In this paper, we reviewed developments in applications of the factorization method to the theory of point vortex patterns that have occurred over the past several decades, and presented new results. These developments are linked to the theory of hierarchies of integrable PDEs and the bi-spectral problem.
Connection with the KdV hierarchy has been known for a long time. The Lax operator of this hierarchy is a Schrodinger operator, and related non-terminating configurations (i.e., $\Lambda=1$ non terminating sequences), both static and translating, on the plane and on the cylinder, can be generated through the factorization of rational or trigonometric Lax-Schrodinger operators. Multi-parametric terminating static configurations, both on the plane and on the cylinder, can also be generated through the factorization of the Schrodinger operator in the $\Lambda=1$ case.
In the $\Lambda=2$ case, connections to the Sawada-Kotera and Kaup-Kupersmidt hierarchies were first found by Demina and Kudryashov [20]. The Lax operators of these hierarchies are of third order, so a natural idea was to consider the Darboux transformations for the third-order operators $L=\partial_{z}^{3}-u\partial_{z}$ [35]. The development of this idea was reviewed in Sections 8 and 9 (also see Appendix 1). The non-terminating and terminating sequences of static configurations on the plane can be generated by such transformations in the $\Lambda=2$ case.
In contrast to the $\Lambda=1$ case, the translating configurations cannot be constructed either through factorization or the $\tau$-function method when $\Lambda=2$. Indeed, in the $\Lambda=1$ casde, the translating configurations are determined by the Baker-Akhiezer function, which is obtained through the action of the intertwining operator on eigenfunctions of the free Lax-Schrodinger operator. However, the Lax operator is of third order in the $\Lambda=2$ (Sawada-Kotera) case. The corresponding third-order equation for the Baker-Akhiezer function of the hierarchy is (see Appendix 1 for details):
$L\Psi(k,z)=k^{3}\Psi(k,z),\quad L=\partial_{z}^{3}+6(\log q(z))^{\prime\prime}\partial_{z},\quad\Psi(k,z)=\frac{\theta(k,z)}{q(z)}e^{kz}.$
It cannot, in general, be reduced to the Schrodinger equation
$H\phi(k,z)=-k^{2}\phi(k,z),\quad H=-\partial_{z}^{2}-\Lambda(\Lambda+1)(\log q(k,z))^{\prime\prime},\quad\phi(k,z)=\frac{p(k,z)}{q(k,z)^{\Lambda}}e^{kz},$
which is equivalent to the bilinear equation for translating configurations. Only when $k=0$ does the eigenvalue problem $L\psi=(\partial_{z}^{2}-u)\partial_{z}\psi=k^{3}\psi$ reduce to solving the Schrodinger equation $(-\partial_{z}^{2}+u)\phi=0$, where $\phi$ is the $k\to 0$ limit of $\psi^{\prime}$:
$\phi(z)=\lim_{k\to 0}\frac{\partial\psi(k,z)}{\partial z},\quad\psi=C_{1}(k)\Psi(k,z)+C_{2}(k)\Psi(e^{2{\rm i}\pi/3}k,z)+C_{3}(k)\Psi(e^{4{\rm i}\pi/3}k,z),$
which corresponds to static configurations (Here, suitably chosen $C_{i}(k)$ ensure finiteness of the limit). As a consequence, translating configurations cannot be generated through Darboux transformations, and one cannot expect to find non-terminating sequences. Thus, the classification of translating configurations remains an open question for $\Lambda=2$.
A similar problem arises when one tries to classify vortex-street patterns in the $\Lambda=2$ case: We recall that non-terminating vortex street configurations in the $\Lambda=1$ case (Section 12) are generated through a chain of Darboux transformations at distinct non-zero levels. The corresponding eigenfunctions result from the action of the trigonometric intertwining operators on the trigonometric eigenfunctions $\sin(k_{i}z+\zeta_{i})$, $k_{i}>0$ of the free Schrodinger operator $H_{0}=-\partial_{z}^{2}$. However, in the $\Lambda=2$ case, we have intertwining between third-order operators, rather than between Schrodinger operators. Moreover, the free Lax operator $L_{0}=\partial_{z}^{3}$ does not have trigonometric eigenfunctions. Therefore, one cannot expect to find non-terminating sequences here.
@@
We searched for sequences of configurations on the cylinder by solving the trigonometric analog of the bilinear equation (10):
$\tau_{n-1}^{\prime\prime}\tau_{n}-2\gamma_{n}\tau_{n-1}^{\prime}\tau_{n}^{\prime}+\gamma_{n}^{2}\tau_{n-1}\tau_{n}^{\prime\prime}+(d_{n-1}-\gamma_{n}d_{n})^{2}\tau_{n-1}\tau_{n}=0,\quad\gamma_{n}\gamma_{n+1}=1,\quad d_{n}=\deg\tau_{n}.$ (122)
@@
In the $\Lambda=2$ case, we found only short terminating sequences. For example
$\tau_{0}=\xi-\frac{1}{\xi},\tau_{1}=\xi^{6}-4\xi^{4}+5\xi^{2}+\frac{5s_{1}}{\xi^{2}}-\frac{4s_{1}}{\xi^{4}}+\frac{s_{1}}{\xi^{6}},\,\tau_{2}=\xi^{5}-5\xi^{3}-s_{2}\xi+\frac{s_{2}}{\xi}-\frac{5s_{1}}{\xi^{3}}+\frac{s_{1}}{\xi^{5}},$
@@
where $\xi=e^{{\rm i}z}$ and $\gamma_{0}=2$. The above example can be continued for one more step if we impose condition $s_{1}=s_{2}^{2}/5$. This gives another terminating sequence with $\tau_{1}=\tau_{1}(\xi,s_{2})$, $\tau_{2}=\tau_{2}(\xi,s_{2})$ and $\tau_{3}=\tau_{3}(\xi,s_{2},s_{3})$ , where $\deg\tau_{3}=14$.
@@
Without going into much detail, we note that, in such sequences, $\tau_{1}$ is a solution to an ODE that can be reduced to a Gauss hypergeometric form [^33] Alternatively, $\tau_{1}$ can be obtained through Darboux transformations of the free Schrodinger operator, so that $\tau_{1}=C_{n}{\cal W}[\sin z,\sin(2z),\sin\left(nz+\zeta\right)]/\sin z$, $C_{n}=-2{\rm i}\exp(-{\rm i}\zeta)/\left((n-1)(n-2)\right)$, and $\deg\tau_{1}=n+2$, $s_{1}=\exp(-2{\rm i}\zeta)$.. Then we can apply KWCC transformations (64) to $\psi_{i}=\tau_{i}^{\gamma_{i}}/\tau_{i-1}$, $i=1,2$, to continue the sequence. This transformation can be applied at most twice. The terminating configurations obtained in this way can depend on no more than two non-trivial free parameters (they depend on two parameters when $\deg\tau_{1}=6,8,10,\dots$).
Thus, the problem of complete classification of vortex patterns on the cylinder in the $\Lambda=2$ case remains open.
As discussed in Sections 4, 13, in the $\Lambda=1$ case, both main and terminating sequences can be expressed in the form of Wronskians. In the $\Lambda=2$ case, configurations of main sequences have Pfaffian representation (see Appendix 1). We note that a determinant representation also exists for the $\Lambda=2$ main sequences, due to connections with the Kaup-Kupershmidt hierarchy (see Section 7 of [35] for more details). Accordingly, the natural question arises about Pfaffian/determinant representations for the $\Lambda=2$ terminating sequences. To address this question, one would need to obtain the Phaffian/determinant representation using the factorization method, rather than the $\tau$-function approach.
@@
Concluding this review, we would like to mention a related problem from the theory of algebraically integrable systems, as our interest in vortex patterns stems from that theory: In the case of the odd family of periodic configurations, it is convenient to rewrite the bilinear chain (121) in terms of homogeneous polynomials in two variables, replacing $\tau_{n}(z)=\prod_{i=1}^{d_{n}}\sin(z-z_{i})$ with
@@
In the new variables, the chain (121) takes the following form: